OpenStudy (anonymous):

The expression for the nth term of a series is (n+2) work out the 3rd and 4th terms of this series

OpenStudy (anonymous):

3rd =5 & 4th =6

OpenStudy (anonymous):

a(n)=n+2 a(3)=3+2=5 a(4)=4+2=6